# Primary Ionization and Particle Identification with Straw Tube Detectors

### v2

@@ 3.2 XTR Study with Straw Tube Detectors
The XTR photons are released when a relativistic particle traverses in an in-homogeneous medium i.e., through the boundary between materials of different electrical properties [28][32]. The relation for transition radiation (TR) production and its dependence on Lorentz factor ($\gamma$) is given by eq. (3.1):
$\frac{dW}{d\omega}=\frac{4\alpha}{\sigma(\kappa+1)}(1-exp(-N_{f}\sigma))\times\Sigma_{n}\theta_{n}\biggl{(}\frac{1}{\rho_{1}+\theta_{n}}-\frac{1}{\rho_{2}+\theta_{n}}\biggr{)}^{2}\biggl{[}1-cos(\rho_{1}+\theta_{n})\biggr{]}$ (3.1)
@@
The eq. (3.1) shows the TR yield and energy spectrum dependence on various factors like Lorentz factor, number of foils, foil thickness and spacing. In the equation, W is energy spectrum radiated by a charged particle. Also, $\alpha$ = 1/137 is the fine structure constant, $\sigma$ is the absorption cross section for the foil and gas. $N_{f}$ represents the number of foils, $\omega$ is plasma frequency for the two media, $\xi_{i}$ <<1, $l_{1}$ is foil thickness, $l_{2}$ is the spacing between the foils, $\theta$ is emission angle. $\beta$ = $v/c$, v is the velocity of particle and c is speed of light. The XTR photon yield gives the amount of XTR photon emitted from the transition of particles (electrons) in different media.
@@ 3.5 Particle Identification with Straw Tube Detectors
The next particle chosen to be identified were kaons. 10,000 kaons were shot in the $XeCO_{2}$ :: 70:30 gas mixture. The physics lists were chosen to be same as that of pions. The kaons and its decay products have been identified in the detector as shown in figure 20. Kaon decays into its several products as shown in the eq. (3.2). These particles deposits their energies in the straw tube detector that has been shown in figure 20. The kaon decays into $\mu^{-}$ and $\pi^{-}$ mainly, $\mu^{-}$ further decays into $e^{-}$ that has been shown in the figure.
$K^{-}\rightarrow\mu^{-}+\bar{\nu}_{\mu}$ (3.2)
@@ 4 Summary and Conclusions
Acknowledgments
I would like to thank the department of physics, Chandigarh University for help and support. I thank the referees for the careful reading of the paper and their suggestions have been valuable in improving the quality of the paper.

### v3

@@ 3.2 XTR Study with Straw Tube Detectors
The XTR photons are released when a relativistic particle traverses in an in-homogeneous medium i.e., through the boundary between materials of different electrical properties [28][32]. The relation for transition radiation (TR) production and its dependence on Lorentz factor ($\gamma$) is given by eq. (1):
$\frac{dW}{d\omega}=\frac{4\alpha}{\sigma(\kappa+1)}(1-exp(-N_{f}\sigma))\times\Sigma_{n}\theta_{n}\biggl{(}\frac{1}{\rho_{1}+\theta_{n}}-\frac{1}{\rho_{2}+\theta_{n}}\biggr{)}^{2}\biggl{[}1-cos(\rho_{1}+\theta_{n})\biggr{]}$ (1)
@@
The eq. (1) shows the TR yield and energy spectrum dependence on various factors like Lorentz factor, number of foils, foil thickness and spacing. In the equation, W is energy spectrum radiated by a charged particle. Also, $\alpha$ = 1/137 is the fine structure constant, $\sigma$ is the absorption cross section for the foil and gas. $N_{f}$ represents the number of foils, $\omega$ is plasma frequency for the two media, $\xi_{i}$¡1, $l_{1}$ is foil thickness, $l_{2}$ is the spacing between the foils, $\theta$ is emission angle. $\beta$ = $v/c$, v is the velocity of particle and c is speed of light. The XTR photon yield gives the amount of XTR photon emitted from the transition of particles (electrons) in different media.
@@ 3.5 Particle Identification with Straw Tube Detectors
The next particle chosen to be identified were kaons. 10,000 kaons were shot in the $XeCO_{2}$ :: 70:30 gas mixture. The physics lists were chosen to be same as that of pions. The kaons and its decay products have been identified in the detector as shown in figure 20. Kaon decays into its several products as shown in the eq. (2). These particles deposits their energies in the straw tube detector that has been shown in figure 20. The kaon decays into $\mu^{-}$ and $\pi^{-}$ mainly, $\mu^{-}$ further decays into $e^{-}$ that has been shown in the figure.
$K^{-}\rightarrow\mu^{-}+\bar{\nu}_{\mu}$ (2)
@@ 4 Summary and Conclusions
Acknowledgements
: I would like to thank the department of physics, Chandigarh University for help and support.